Iron can be obtained by reacting 200g of Fe2O3 with 42g of CO.

Fe2O3 + 3CO → 2Fe + 3CO2

How much of the excess reagent in grams is left at the end of the reaction? And, what will be the percent yield if 48g of Fe is the actual yield?


1
Expert's answer
2021-10-17T11:18:26-0400

Fe2O3 + 3CO → 2Fe + 3CO2


Mole ratio:-1:3:2:3

Moles of iron yielded "=\frac{mass}{RAM}"

"=\frac{48g}{56g}=0.8571moles"


Mole ratio of Fe2O3:Fe from the equation above is 1:2

"\therefore" Moles of Fe2O3 reaction="\frac{1}{2}" ×Moles of Fe yielded.

="\frac{1}{2}×0.8571moles" =0.42855moles


But moles of Fe2O3 reacting ="\frac{mass reacting}{RFM of Fe_2 O_3}"


"0.42855=\frac{mass}{(56×2)+(16×3)}"


"\therefore" Mass of Fe2O3 reacted =0.42855×160=68.568g

But mass of excess Fe2O3 =200g -68.568g =131.432g

Mass of excess Fe2O3=131.432g


%yield if 48g of Fe is the actual yield="\frac{68.568g}{200g}×" 100%=34.284%


=34.284%